Y=5x-3 with y=2x find the solution of each system of simultaneous equations by graphing
Answer by mananth(16949)

-5 x + 1 y = -3 .............1
-2 x 1 y = 0 .............2
Find the x & y intercept points by plugging x=0,y=0 in both equations.
equation I
when x= 0 y= -3 ,( 0 , -3 )
when y = 0 x= -3 ,( 0.6 , 0 )
Equation 2
when x= 0 y= 0 ( 0 , 0 )
when y = 0 x= 0 ( 0 , 0 )



You will observe that the point of intersection is 1,2
The solution of the system is ( 1,2 )